# Refinable modeling for unbinned SMEFT analyses

### v1

@@ 4 Simulation for inference
Despite the computational cost, the robust and reliable predictions from Monte Carlo (MC) simulations are a core strength in modeling highly energetic scattering processes [29]. In this section, we extend the methods for modeling systematic effects to the unbinned case, enabling us to leverage the extensive experience from over a decade of binned LHC data analyses. Instead of using simulated variations in binned cumulative yields, we train ML surrogates on the simulations to provide continuous estimates in $\boldsymbol{x}$ and ${\boldsymbol{\nu}}$. Given the diverse nature of existing systematic modeling procedures, a detailed statistical interpretation of event simulation is necessary.
@@ 4.1 Hierarchical data representations and staged event simulation
Systematic effects, in general, can modify the predictions at any energy scale. As shown in Fig. 1, we broadly categorize accordingly in the parton level (p), the particle level (ptl), the detector reconstruction level (reco), and observed features, derived from the reconstruction level and denoted by $\boldsymbol{x}$. By definition, the observables $\boldsymbol{x}$ are the quantities entering Eqs. 4 - 11. All other quantities are denoted by $\boldsymbol{z}$, including $\boldsymbol{z}_{\textrm{reco}}$, which contains, for example, low-level hit patterns in sub-detectors that can, in principle, be accessed in real data.
@@
This coarse grouping is a compromise, balancing the detail level with the notational clutter. It also reflects the approximate separation of systematic effects at the UV and SMEFT energy scales from the QCD and detector signals’ energy scales. This staging is not a constraint; extra stages can be introduced as long as an event representation can be defined.
@@
Technically, event simulation for LHC proton-proton (pp) collisions is split into several computer codes that address the modeling challenge at each energy scale with different techniques. At the parton level, ME event generators such as MG5_aMC@NLO [30, 31], sherpa [32], or powheg [33, 34, 35, 36, 37] provide a sampling of purely perturbative ME-squared predictions for the hard-scatter interaction. We represent the dynamics of fundamental particles emanating from the hard-scatter and the subsequent decays of the heavy SM particles by $\boldsymbol{z}_{\textrm{p}}$. The PS evolves $\boldsymbol{z}_{\textrm{p}}$ down to energies where the perturbative approach is no longer viable. With the simulation of the hadronization, color reconnection, the underlying event, the decay of unstable hadrons, the particles from multiple-parton interactions, and the initial-state and final-state radiation, the PS belongs to the particle level. It is simulated with general-purpose software tools such as pythia [38] or herwig [39] and requires tuning of the involved phenomenological parameters to describe the data reliably. The resulting particle-level event representation is denoted by $\boldsymbol{z}_{\textrm{ptl}}$. The parton and particle levels are not available in real data and are always latent.
@@
Particle-level events are further processed with detector-specific simulation such as geant [40], using the conditions appropriate for each data-taking period and mixed with the simulation of separate hard-scatter events to model pile-up. After the stage of simulated detector hits, the event reconstruction proceeds up to the particle candidate level, for example, provided by a version of the ParticleFlow algorithm [41, 42]. Together with the jet clustering, the identification and disambiguation of leptons, etc., it proceeds similarly in real and simulated data. The result constitutes the reconstruction level event whose features we denote by $\boldsymbol{z}_{\textrm{reco}}$.
@@
Most data analyses derive high-level observables $\boldsymbol{x}$ from $\boldsymbol{z}_{\textrm{reco}}$. These observables represent all event features we wish to include in the hypothesis test. In principle, $\boldsymbol{x}$ can encompass the entire reconstruction level, including the variable-length list of all reconstructed particle candidates in an event [43]. Such approaches have even been used to constrain SMEFT effects [44]; here, we restrict to the case of high-level event features derived from the particle candidates. Simplified event reconstruction is available from, e.g., Delphes [45].
@@
For bookkeeping, we similarly group the nuisance parameters into ${\boldsymbol{\nu}}_{\textrm{p}}$, ${\boldsymbol{\nu}}_{\textrm{ptl}}$, and ${\boldsymbol{\nu}}_{\textrm{reco}}$ which we collectively denote by ${\boldsymbol{\nu}}$ when it simplifies the notation. The differential cross-section in the fiducial phase space in Eq. 4 becomes
@@
The main advantage of the hierarchical event representations is a well-motivated factorization of the pdf according to
@@
This pdf jointly depends on latent and observable features [14]. In ratios, the various conditional factors [^1] The factor $p(\boldsymbol{x}|\boldsymbol{z}_{\textrm{reco}})$ could also be conditional on nuisance parameters related to uncertainties in analysis-dependent parameters that may be involved when computing $\boldsymbol{x}$ from $\boldsymbol{z}_{\textrm{reco}}$. This extension is straightforward. can partially or entirely cancel, enabling efficient generation of synthetic data sets that, in turn, facilitate the training of ML surrogates [17].
@@ 4.2 Semi-analytic modeling at the parton level
Uncertainties in the PDFs can, for example, be given in terms of variations in Hessian eigen-directions of an underlying parametrization [46] and we denote the corresponding nuisance parameters by ${\boldsymbol{\nu}}_{\textrm{PDF}}$. The PDF variations describe different hypotheses on the parton-dynamics of the proton and thus have a meaningful physical interpretation. The same can not be said for the dependence on $\mu_{R}$ and $\mu_{F}$, which are technical artifacts of the finite perturbative order and the separation of collinear radiation between the ME and the PDF. With this caveat in mind, the scale uncertainties are nevertheless treated with standard nuisance parameters $\nu_{R}$ and $\nu_{F}$ corresponding to up and down variations around the central values $\mu_{R,0}$ and $\mu_{R,0}$. If we take values of ±$\pm 1$ to correspond to variations factors of 2, we have
@@
We account for uncertainties in the overall normalization of the process with log-normal nuisances in Sec. 6 and do not discuss these here. However, we still include a more general parton-level ad-hoc reweighting function $\alpha_{\textrm{rw}}(\boldsymbol{z}_{\textrm{p}},{\boldsymbol{\nu}}_{\textrm{rw}})$. It can be employed, for example, if higher-order perturbative corrections significantly depend on the parton-level configuration and we wish to modify the parton-level distribution with an ad-hoc modification. We thus write the total parton-level prediction as
@@
An example of this type is the modeling of the transverse top quark momentum in the $\textrm{t}\bar{\textrm{t}}$ process, thought to be well parametrized by the parton-level transverse top momenta $p_{\textrm{T}}(\textrm{t})$ and $p_{\textrm{T}}(\overline{\textrm{t}})$ [48] but which so far can not be provided by SMEFT ME generators. The nuisances ${\boldsymbol{\nu}}_{\textrm{rw}}$ then modify parameters in the reweighting function $\alpha_{\textrm{rw}}(\boldsymbol{z}_{p},{\boldsymbol{\nu}}_{\textrm{rw}})$ within their uncertainties. Because such procedures are highly application-dependent, we do not go into further detail except to note that $\alpha_{\textrm{rw}}(\boldsymbol{z}_{p},{\boldsymbol{\nu}}_{\textrm{rw}})$ must always be positive. The parton-level nuisance parameters discussed so far are
@@
and they are related to systematic effects that we can model semi-analytically, i.e., by an efficient computation.
@@
The bar on $\bar{\sigma}({{\boldsymbol{\theta}}},{\boldsymbol{\nu}}_{\textrm{p}})$ indicates that the inclusive cross section pertains to the entire kinematic phase-space, unaffected, e.g., by the finite detector acceptance. By definition, we have
@@
These parton-level systematic effects allow tractable simulation, i.e., there is an inexpensive way of modifying an existing sample, for example, by re-weighting, so that it provides an approximation for model parameters other than its nominal ones. Many other effects, for example, the choice among different ME generators, do not allow tractable simulation. If differences of such “two-point alternatives” are modeled with nuisance parameters, we can not compute the likelihood ratio corresponding to one generator choice when sampling from another. Nevertheless, we show in Sec. 5 how to machine-learn parametrizations that interpolate between two-point modeling alternatives using a nuisance parameter. This detail is of practical importance because uncertainties, for example, in modeling non-perturbative aspects of QCD, are frequently of this type. Two-point alternatives are, therefore, also encountered at the particle and detector level.
@@ 4.3 Forward-mode event generation at particle and reconstruction level
Particle-level simulation, staged as described in Sec. 4.1, simulates the PS, initial-state and final-state radiation, multiple-parton interactions, color reconnection, hadronization, the underlying event, and hadron decay. The subsequent reconstruction-level simulation models the detector interaction and event reconstruction. For every parton-level event $\{w_{i},\boldsymbol{z}_{\textrm{p},i}\}$, the particle and detector-level simulators and the reconstruction of the observables provide an event representation of the form
@@
Because of the finite detector acceptance, some events will fail the event selection. The typical poor reconstruction performance close to reconstruction thresholds motivates the definition of a fiducial region, which we denote by $\mathcal{X}$. The need to reduce backgrounds, the availability of reliable object-level calibrations, and apriori knowledge about SMEFT-sensitivity can all be addressed by analysis-specific selections that we incorporate in the definition of $\mathcal{X}$. We only require that for a real event, it must be possible to decide whether or not it is included in $\mathcal{X}$. In particular, $\mathcal{X}$ can include requirements on the latent $\boldsymbol{z}_{\textrm{reco}}$: acceptance at the online selection stage, thresholds on properties of the reconstructed objects, and the various event vetos for data-cleaning are all part of the definition of $\mathcal{X}$ despite these variables are usually not in $\boldsymbol{x}$.
@@ 4.4 Synthetic data sets and tractable simulation
Simulated samples are computationally expensive. While it is always possible to obtain a simulation according to Eq. 23 for a specific model parameter configuration $({\boldsymbol{\theta}_{0}},{{\boldsymbol{\nu}}_{0}})$, it makes a significant practical difference whether or not we can efficiently produce a new simulated (synthetic) data set from an existing one if ${{\boldsymbol{\theta}}}\neq{\boldsymbol{\theta}_{0}}$ or ${\boldsymbol{\nu}}\neq{{\boldsymbol{\nu}}_{0}}$ for at least some model parameters. The simulation is then called “tractable” for these parameters. We discuss two such tractable cases: likelihood-based reweighting and variations in the calibration of the underlying reconstructed objects. Because we can also model the SMEFT effects tractably, we separately discuss this case. If a simulation is non-tractable for a model parameter, we must resort to Eq. 23 to obtain systematically varied data sets. A visualization of the various approaches is provided in Fig. 2.
@@ 4.4.1 Uncertainties in the calibration of reconstructed objects
An essential class of tractable simulation pertains to effects from uncertainties in the calibration of the underlying objects’ properties, such as jet and lepton momenta or the discriminator value of a $\mathrm{b}$-tagging algorithm, etc. The corresponding variations of $\boldsymbol{x}$ can be obtained by recomputing $\boldsymbol{x}$ from varied event properties, defining a function $\boldsymbol{x}_{\boldsymbol{\nu}}=J_{{\boldsymbol{\nu}}}(\boldsymbol{x},\boldsymbol{z})$. It provides modified values of $\boldsymbol{x}$ that can depend on latent object-level and event properties. Hence, $J_{{\boldsymbol{\nu}}}$ also depends on $\boldsymbol{z}$.
@@
i.e., the joint likelihood does not change with ${\boldsymbol{\nu}}$ when we simultaneously modify the observation to $\boldsymbol{x}_{{\boldsymbol{\nu}},i}=J_{{\boldsymbol{\nu}}}(\boldsymbol{x}_{i},\boldsymbol{z}_{i})$. Synthetic data samples are efficiently obtained as
@@
and are enough to learn a DCR surrogate for two values of ${\boldsymbol{\nu}}$ for the same observation, as we will discuss in Sec. 5. We have assumed the independence of calibration-type uncertainties on the POIs: the same value ${\boldsymbol{\theta}_{0}}$ appears on both sides of Eq. 29
@@ 4.4.2 Synthetic data from event-reweighting
When the change in the likelihood of observing an event as a function of a specific model parameter can be computed without resampling the pdf, we can obtain reweighted synthetic data sets as
@@
A reweighting function $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}},{\boldsymbol{\theta}_{0}},{{\boldsymbol{\nu}}_{0}})$ must be available for the corresponding model parameter. It can depend on both observables and latent features. Important tractable examples pertain to the parton level, where access to the analytic implementation of the ME-squared terms allows the computation of the joint likelihood ratio. Besides parton-level parameterizations of the form
@@
that are designed to mimic the effect of higher-order perturbative corrections as in Eq. 17, it can be useful to describe the uncertainties related to other inaccuracies of the theoretical predictions with nuisance parameters. For example, the predicted rates of events with a high multiplicity of generated particle-level jets ($N_{\textrm{gen-jet}}$) depend on the details of the matching of the ME calculation and the PS [50], with significant uncertainties in many cases. We can, for example, associate ad-hoc uncertainties to events with different particle-level jet multiplicities using the ad-hoc variation
@@
with constants $\alpha_{1}$, $\alpha_{2}$, etc. and a corresponding nuisance parameter $\nu_{\textrm{gen-jet}}$.
A particular case is reweighting functions that solely depend on $\boldsymbol{x}$. In this case, we can access the DCR as a function of the observed features, and nothing must be learned. As an example, consider a selection with a fixed lepton multiplicity $N_{\ell}$ whose reconstruction efficiency is subject to a relative uncertainty $\Delta\textrm{SF}(\ell)/\textrm{SF}(\ell)$ where a corrective scale factor $\textrm{SF}(\ell)$ is already accounted for in the nominal simulation. We wish to treat its uncertainty $\Delta\textrm{SF}(\ell)$ with a nuisance parameter $\nu_{\ell}$. If $\boldsymbol{x}$ encompasses the properties of the leptons and $\Delta\textrm{SF}(\ell)$ can thus be computed solely from $\boldsymbol{x}$ we have
$r(\boldsymbol{x}_{i}|\nu_{\ell})=\prod_{i=1}^{N_{\ell}}\left(1+\frac{\Delta\textrm{SF}(\ell_{i})}{\textrm{SF}(\ell_{i})}\right)^{{\boldsymbol{\nu}}_{\ell}},$ (34)
which, in this case, provides the DCR as a function of $\nu_{\ell}$. No surrogate is needed. If $\boldsymbol{x}$ is not sufficient for computing the scale factors, the r.h.s. of Eq. 34 does not equate to $r(\boldsymbol{x}_{i}|\nu)$ but $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|\nu)$ and a surrogate is required.

### v3

@@ 4 Simulation for inference
Although computationally intensive, robust predictions from Monte Carlo (MC) simulations are invaluable for modeling highly energetic scattering processes [30]. Our strategy is to provide the necessary parametric estimates by breaking down the primary task of modeling the DCR into smaller, manageable machine-learning tasks. The tasks are based on simulation and capitalize on the high quality of the MC methods. As illustrated in Figure 2, we proceed in the backward mode, from left to right, and separate the training into distinct parton-level processes that are simulated separately. For each process, we learn parametrizations of systematic effects and POI dependencies with the help of efficiently generated “synthetic” data sets that correspond to the systematic variations in the binned approach. Inheriting these procedures, we leverage the extensive experience from over a decade of binned LHC data analyses.
Training tasks for systematics can be divided into uncorrelated groups of nuisance parameters which then separately estimate these effects. A high granularity allows a gradual refinement of each aspect of the final model without invalidating unrelated tasks. In the following sections, we describe how the results of ML training tasks combine into a refinable surrogate model ${\textrm{d}}\Sigma(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ that provides the needed DCR in Eq. 11.
@@ 4.1 Hierarchical data representations and staged event simulation
Systematic effects can modify predictions at any energy scale. As shown in Fig. 1, we broadly categorize these into parton level (p), particle level (ptl), detector reconstruction level (reco), and observed features derived from reconstruction, denoted by $\boldsymbol{x}$. By definition, the observables $\boldsymbol{x}$ are the quantities used in Eqs. 4 - 11. All other quantities, including $\boldsymbol{z}_{\textrm{reco}}$ (e.g., low-level hit patterns in sub-detectors accessible in real data), are denoted by $\boldsymbol{z}$.
@@
This grouping strikes a balance between detail and notational simplicity. It also reflects the approximate separation of systematic effects from UV and SMEFT energy scales to those relevant to QCD and detector signals. This staging is flexible; additional stages can be introduced as long as an event representation can be defined.
@@
Event simulation for LHC proton-proton (pp) collisions is divided among several computer codes, each addressing modeling at a specific energy scale with specialized techniques. At the parton level, ME generators like MG5_aMC@NLO [31, 32], sherpa [33], and powheg [34, 35, 36, 37, 38] provide a sampling of purely perturbative ME-squared predictions for the hard-scatter interaction. The dynamics of fundamental particles from the hard scatter and subsequent decays of heavy SM particles are represented by $\boldsymbol{z}_{\textrm{p}}$. The PS evolves $\boldsymbol{z}_{\textrm{p}}$ down to energies where perturbative methods are no longer valid. Together with color reconnection, hadronization, decays of unstable hadrons, underlying event, particles from multiple-parton interactions, and both initial- and final-state radiation, it defines the particle level. This is simulated with general-purpose tools like pythia [39] or herwig [40], which require tuning of phenomenological parameters to reliably describe data. The resulting particle-level event is represented by $\boldsymbol{z}_{\textrm{ptl}}$. The parton and particle levels are latent and cannot be directly observed in real data.
@@
Particle-level events are processed with detector-specific simulation tools like geant [41], using conditions for each data-taking period and mixed with simulations of separate hard-scatter events to model pile-up. After simulating detector hits, event reconstruction proceeds to the particle candidate level, using, e.g., a version of the ParticleFlow algorithm [42, 43]. Together with jet clustering, lepton identification, and disambiguation, this process is similar in both real and simulated data. The result is the reconstruction-level event, with features denoted by $\boldsymbol{z}_{\textrm{reco}}$. Simplified event reconstruction is available from, e.g., Delphes [44].
@@
Most data analyses derive high-level observables $\boldsymbol{x}$ from $\boldsymbol{z}_{\textrm{reco}}$. These observables capture all event features included in the hypothesis test. In principle, $\boldsymbol{x}$ can represent the entire reconstruction level, including the variable-length list of all reconstructed particle candidates in an event [45]. Such approaches have been used to constrain SMEFT effects [46]; here, we focus on high-level event features. Any real-data event features not already included in $\boldsymbol{x}$ belong to $\boldsymbol{z}_{\textrm{reco}}$ and are latent in the hypothesis test.
@@
For bookkeeping, we group the nuisance parameters into ${\boldsymbol{\nu}}_{\textrm{p}}$, ${\boldsymbol{\nu}}_{\textrm{ptl}}$, and ${\boldsymbol{\nu}}_{\textrm{reco}}$, collectively denoted by ${\boldsymbol{\nu}}$ when simplifying notation. The differential cross-section in the fiducial phase space from Eq. 4 then becomes
@@
The hierarchical event representation enables a natural factorization of the pdf as
@@
This pdf depends on both latent and observable features [14]. In ratios, the conditional factors [^1] The factor $p(\boldsymbol{x}|\boldsymbol{z}_{\textrm{reco}})$ could also be conditional on nuisance parameters related to uncertainties in analysis-dependent parameters that may be involved when computing $\boldsymbol{x}$ from $\boldsymbol{z}_{\textrm{reco}}$. This extension is straightforward. can partially or entirely cancel, enabling efficient generation of synthetic data sets for trainng ML surrogates [17].
@@ 4.2 Semi-analytic modeling at the parton level
Uncertainties in the PDFs can be expressed as variations along Hessian eigen-directions of an underlying parametrization [47], with the corresponding nuisance parameters denoted by ${\boldsymbol{\nu}}_{\textrm{PDF}}$. These PDF variations represent different hypotheses about proton parton dynamics and have a clear physical interpretation. In contrast, the dependence on $\mu_{R}$ and $\mu_{F}$ arises from technical artifacts, namely the finite perturbative order and the separation of collinear radiation between the ME and the PDF. Despite this, we treat scale uncertainties with standard nuisance parameters $\nu_{R}$ and $\nu_{F}$, corresponding to up and down variations around the central values $\mu_{R,0}$ and $\mu_{F,0}$. If values of ±$\pm 1$ correspond to variation factors of 2, then
@@
We account for uncertainties in the overall normalization of the process with log-normal nuisances in Sec. 6 and will not discuss them here. However, we include a general parton-level ad-hoc reweighting function $\alpha_{\textrm{rw}}(\boldsymbol{z}_{\textrm{p}},{\boldsymbol{\nu}}_{\textrm{rw}})$, useful when higher-order perturbative corrections significantly depend on the parton-level configuration and we want to adjust the parton-level distribution with an ad-hoc modification. The total parton-level prediction is then written as
@@
An example of this type is the modeling of the transverse top quark momenta in the $\textrm{t}\bar{\textrm{t}}$ process, which is well understood from higher-order pertubation theory [49], but is not yet available from SMEFT ME generators. The nuisances ${\boldsymbol{\nu}}_{\textrm{rw}}$ modify parameters in the reweighting function $\alpha_{\textrm{rw}}(\boldsymbol{z}_{p},{\boldsymbol{\nu}}_{\textrm{rw}})$ within their uncertainties. Since such procedures are highly application-dependent, we do not elaborate further, except to note that $\alpha_{\textrm{rw}}(\boldsymbol{z}_{p},{\boldsymbol{\nu}}_{\textrm{rw}})$ must always be positive. The parton-level nuisance parameters considered so far are
@@
and are associated with systematic effects that can be modeled semi-analytically, allowing for efficient computation.
@@
The bar on $\bar{\sigma}({{\boldsymbol{\theta}}},{\boldsymbol{\nu}}_{\textrm{p}})$ indicates that the inclusive cross-section pertains to the entire kinematic phase-space, unaffected, e.g., by the finite detector acceptance. By definition, we have
@@
These parton-level systematic effects enable tractable simulation, allowing an inexpensive way to modify an existing sample, such as by reweighting, to approximate model parameters beyond the nominal values. Many other effects, such as the choice of different ME generators, do not allow for tractable simulation. If we model differences between such “two point alternatives” with nuisance parameters, it becomes impossible to compute the likelihood ratio for one generator choice while sampling from another. However, in Sec. 5, we show how to use machine learning to create parametrizations that interpolate between two point modeling alternatives using a nuisance parameter. This approach is crucial in practice, as uncertainties, especially those in modeling nonperturbative aspects of QCD, are often of this type.
@@ 4.3 Forward-mode event generation at particle and reconstruction level
Particle-level simulation, staged as described in Sec. 4.1, includes the PS, initial- and final-state radiation, multiple-parton interactions, color reconnection, hadronization, the underlying event, and hadron decay. The subsequent reconstruction-level simulation models the detector interaction and event reconstruction. For each parton-level event $\{w_{i},\boldsymbol{z}_{\textrm{p},i}\}$, the particle and detector-level simulations, along with observable reconstruction, produce an event representation of the form
@@
Due to finite detector acceptance, some events will not pass the event selection. Poor reconstruction performance near reconstruction thresholds motivates defining a fiducial region, denoted by $\mathcal{X}$. Analysis-specific selections, including reliable object-level calibrations, background reduction, and prior knowledge of SMEFT sensitivity, are incorporated into the definition of $\mathcal{X}$. We only require that for a real event, it is possible to determine if it belongs in $\mathcal{X}$. In particular, $\mathcal{X}$ can include requirements on latent variables like $\boldsymbol{z}_{\textrm{reco}}$: online selection acceptance, thresholds on reconstructed object properties, and various data-cleaning event vetoes are part of the definition of $\mathcal{X}$, even if these variables are typically not in $\boldsymbol{x}$.
@@ 4.4 Synthetic data sets and tractable simulation
Simulated samples are computationally expensive. While it is always possible to obtain a simulation based on Eq. 23 for a specific model parameter configuration $({\boldsymbol{\theta}_{0}},{{\boldsymbol{\nu}}_{0}})$, it is practically important to know whether we can efficiently generate a new simulated (synthetic) data set from an existing one when ${{\boldsymbol{\theta}}}\neq{\boldsymbol{\theta}_{0}}$ or ${\boldsymbol{\nu}}\neq{{\boldsymbol{\nu}}_{0}}$ for some model parameters. In such cases, the simulation is called “tractable” for these parameters.
We discuss two tractable cases: likelihood-based reweighting and variations in the calibration of the reconstructed objects. Since SMEFT effects can also be modeled tractably, we address this separately. If a simulation is non-tractable for a model parameter, we must use Eq. 23 to obtain systematically varied data sets. A visualization of these approaches is shown in Fig. 3.
@@ 4.4.1 Uncertainties in the calibration of reconstructed objects
An important type of tractable simulation addresses uncertainties in the calibration of underlying object properties, such as jet and lepton momenta or the discriminator value of a $\mathrm{b}$-tagging algorithm. Variations of $\boldsymbol{x}$ from these uncertainties are obtained by recomputing it based on modified event properties, defining a function $\boldsymbol{x}_{\boldsymbol{\nu}}=J_{{\boldsymbol{\nu}}}(\boldsymbol{x},\boldsymbol{z})$. This function provides adjusted values of $\boldsymbol{x}$ that depend on latent object-level and event properties, so $J_{{\boldsymbol{\nu}}}$ also depends on $\boldsymbol{z}$.
@@
meaning the joint likelihood remains unchanged with ${\boldsymbol{\nu}}$ when we simultaneously modify the observation to $\boldsymbol{x}_{{\boldsymbol{\nu}},i}=J_{{\boldsymbol{\nu}}}(\boldsymbol{x}_{i},\boldsymbol{z}_{i})$. Synthetic data samples can be efficiently generated as
@@
and are enough to learn a DCR surrogate as a function of ${\boldsymbol{\nu}}$, as discussed in Sec. 5. We assume that calibration-type uncertainties are independent of the POIs, so ${\boldsymbol{\theta}_{0}}$ appears on both sides of Eq. 29.
@@ 4.4.2 Synthetic data from event-reweighting
When the change in the likelihood of observing an event as a function of a specific model parameter can be computed without resampling the pdf, we can generate reweighted synthetic data sets as
@@
A reweighting function $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}},{\boldsymbol{\theta}_{0}},{{\boldsymbol{\nu}}_{0}})$ must be available for the corresponding model parameter and can depend on both observables and latent features. Important tractable cases occur at the parton level, where access to the analytic form of ME-squared terms allows computation of the joint likelihood ratio. In addition to parton-level parametrizations of the form
@@
designed to approximate higher-order perturbative corrections as in Eq. 17, it is useful to describe uncertainties related to other theoretical inaccuracies with nuisance parameters. For example, the predicted rates of events with high particle-level jet multiplicity ($N_{\textrm{gen-jet}}$) depend on the specifics of matching between the ME calculation and the PS [51], often with significant uncertainties. To address this, we can introduce ad-hoc uncertainties for events with different particle-level jet multiplicities through the variation
@@
where $\alpha_{1}$, $\alpha_{2}$, etc., are constants, and $\nu_{\textrm{gen-jet}}$ is the associated nuisance parameter.
A special case is reweighting functions that depend solely on $\boldsymbol{x}$. In this case, we can access the DCR as a function of the observed features without requiring any learning. For example, consider a selection with a fixed lepton multiplicity $N_{\ell}$, where the reconstruction efficiency has a relative uncertainty $\Delta\textrm{SF}(\ell)/\textrm{SF}(\ell)$, and a corrective scale factor $\textrm{SF}(\ell)$ is already included in the nominal simulation. We treat the uncertainty $\Delta\textrm{SF}(\ell)$ with a nuisance parameter $\nu_{\ell}$. If $\boldsymbol{x}$ includes the properties of the leptons, allowing $\Delta\textrm{SF}(\ell)$ to be computed solely from $\boldsymbol{x}$, we have
$r(\boldsymbol{x}_{i}|\nu_{\ell})=\prod_{i=1}^{N_{\ell}}\left(1+\frac{\Delta\textrm{SF}(\ell_{i})}{\textrm{SF}(\ell_{i})}\right)^{\nu_{\ell}},$ (34)
which provides the DCR as a function of $\nu_{\ell}$ without needing a surrogate. However, if $\boldsymbol{x}$ alone is insufficient to compute the scale factors, the right-hand side of Eq. 34 represents $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|\nu)$ rather than $r(\boldsymbol{x}_{i}|\nu)$, and a surrogate is required.